OpenStudy (mathmath333):

\(\large \color{black}{\begin{align}& \normalsize \text{Find the domain of the function} \hspace{.33em}\\~\\ & f(x)=\log_{10} [1-\log_{10} (x^2-5x+16)] \hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (anonymous):

First thing's first - do you know what the domain of log (a) is ?

OpenStudy (mathmath333):

a>0

OpenStudy (anonymous):

Precisely! So from that we can deduce so much that 1 - log (x^2-5x+16) has to be >0 for that whole thing to exist.

OpenStudy (anonymous):

From that we can deduce that log (x^2-5x+16) has to be <1 so that 1 - log (x^2-5x+16) >0

OpenStudy (mathmath333):

ok

OpenStudy (anonymous):

But at the same time, for log (x^2-5x+16) to exist, (x^2-5x+16) also has to be >0.

OpenStudy (michele_laino):

we have this condition, for the existence of the outer logarithm: \[\Large \begin{gathered} 1 - {\log _{10}}\left( {{x^2} - 5x + 16} \right) > 0 \hfill \\ \hfill \\ {\log _{10}}\left[ {\frac{{10}}{{{x^2} - 5x + 16}}} \right] > 0 \hfill \\ \hfill \\ \frac{{10}}{{{x^2} - 5x + 16}} > 1 \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

furthermore we have the subsequent condition for the existence of the inner logarithm: \[\Large {x^2} - 5x + 16 > 0\]

OpenStudy (michele_laino):

then the domain of our function is given by the solution of this system: \[\Large \left\{ \begin{gathered} \frac{{10}}{{{x^2} - 5x + 16}} > 1 \hfill \\ \hfill \\ {x^2} - 5x + 16 > 0 \hfill \\ \end{gathered} \right.\]

OpenStudy (anonymous):

It sounds a little wrapped up but it will make sense along the way. Let's work this one step at a time. For log (x^2-5x+16) to exist, (x^2-5x+16) also has to be >0. We solve the quadratic we find that (x^2-5x+16) has complex roots and that means that (x^2-5x+16) is always positive so we need not worry about it! Now, for that whole thing to exist, we've concluded that 1 - log (x^2-5x+16) >0 We "flip" the log over a little and we have that 1 > log (x^2-5x+16) Express 1 as log (10) ( log(10) is 1 ) and you have log (10) > log (x^2-5x+16) Which equates to 10 > (x^2-5x+16) Or 0 > (x^2-5x+6) We solve this new quadratic and we have that (x^2-5x+6)=(x-2)*(x-3). What that means is that (x^2-5x+6) will look something like this: